# Intuitive Axial Augmentation Using Polar-Sine-Based Piecewise Distortion for Medical Slice-Wise Segmentation

### v1

@@ 4.2 Similarity Guide Results
We observe that the number of successful SIFT pairings starts to plummet with $a\approx 3$$f\approx 1.5$ fig. 9. The neural network ablation experiments also exhibit an exciting parallel trend: accuracy starts to falter rather than improve when the augmentation intensity with $A>7,\omega>3$. The Pearson correlation coefficient between model performance and two similarity metrics are $\rho_{SIRF-Acc}=-0.625$ and $\rho_{ORB-Acc}=-0.618$.
This proves that we can use similarity detection to quickly and effectively predict the performance of neural networks under this parameter without any actual training.
(a) SIFT
(b) ORB
(c) Segment Acc.
(d) Correlation
Figure 9: Similarity guide results. Relations between the Count of Feature Points Matched Successfully and Actual Segmentation Performance. The feature points are extracted by SIRF and ORB. The Z-axis uses a logarithmic scale to more stably observe the level of correlation. When $\omega$ exceeds $2$, a significant decline in correlation is observed; in contrast, the effect of $A$ on the correlation is relatively mild. The third subfigure also shows the same trend (red arrow). The far right subfigure is the Pearson correlation coefficient heatmap. A significant negative correlation is observed between similarity and the achievable accuracy.
Given that our annotations are focused on gastric lesions, we have conducted an in-depth examination of the similarity levels across upper abdomen. This is aimed at assisting in a more comprehensive determination of the optimal distortion intensity fig. 10. The similarity criterion is described in eq. 13. A higher $N_{\triangle}$ indicates a greater complexity of textures in the respective axial position, enabling the extraction of a larger number of feature points.
@@
(a) Similarity Metric
(b) Slice Gap on Coronal Section
Figure 10: Similarity distribution around labelled slices. The vertical axis in the first and second subfigure correspond to SIFT and ORB value. The number of similarity detection points shows an upward trend when $\triangle\rightarrow u,u\notin stomach\approx\left[-20,20\right]$. The similarity results indicate a scarcity of distinctive features in the axial slice where the stomach is located. This aligns with anatomical realities: the heart and lungs are characterized by a higher density of blood vessels and bronchi, while the mid-to-lower abdomen contains a diverse array of organs.
4.3 Segmentation Performances
We examine our method’s effectiveness the latest segmentation models, MedNext[26] and SwinUMamba[14], combined with the large dataset provided by SA-Med2D[33]. The results are presented in table 4.
In table 6, we evaluated our method on several traditional and widely-used neural networks, training on our private dataset. We improve segmentation accuracy across multiple neural-network-based segmentation frameworks, which are selected from the most representative ones in recent computer vision researches. Most frameworks can steady gain higher value on major metrics without the extra samples. The straightforward improvement show that our method is robust across different designs of neural networks. fig. 11 illustrate the Precision-recall Curve. Due to limited computational resources, we focused on calculating the most dynamically changing segment of the PR curve.
Table 4: Performance on the large CT-ORG[25] dataset. Our method is able to help the latest neural networks to achieve higher accuracy with limited training samples, specially useful in medical scenarios. This dataset is a typical representative of CT modality.
@@ 4.3 Segmentation Performances
MedNext Vanilla Dice 96.39 65.53 3.21 19.32 66.26 86.13 56.14
@@
SwinUMamba Vanilla Dice 96.85 71.44 1.39 21.38 71.84 92.3 59.21
@@
Table 5: The similar experiment with table 4, but was on the BraTS[20] dataset, and trained with SwinUMamba. This dataset is a typical representative of MR modality.
@@
BraTS 2020 Vanilla Dice 99.06 37.69 13.75 26.93 44.36
@@
BraTS 2021 Vanilla Dice 99.03 35.24 1.53 30.76 41.64
@@
Family Model Criterion w/o aug1 rotate ±$\pm 45^{\circ}$ rotate ±$\pm 90^{\circ}$ rotate ±$\pm 180^{\circ}$ Ours
Conv-Based Resnet50 mIoU 72.52 78.32 79.49 78.58 82.57
mDice 81.28 86.28 87.20 86.48 89.52
mRecall 82.39 87.34 88.67 91.53 91.37
mPrecision 83.81 87.04 87.80 85.87 89.05
ConvNext mIoU 63.47 71.99 75.2 75.00 79.64
mDice 71.72 80.77 83.75 83.51 87.32
mRecall 77.36 82.26 84.36 86.95 90.37
mPrecision 75.54 82.57 83.75 82.83 87.16
SegNeXt mIoU 64.02 72.46 73.04 74.16 76.61
mDice 72.37 81.23 81.77 82.76 84.88
mRecall 70.31 83.81 86.20 85.98 89.39
mPrecision 77.46 80.73 82.54 83.60 85.57
Trans-Based MAE mIoU 67.81 71.73 70.00 68.90 75.04
mDice 76.63 80.56 78.89 77.81 83.55
mRecall 85.79 85.80 82.25 84.59 87.09
mPrecision 78.27 77.44 78.66 73.90 82.70
Poolformer mIoU 68.08 76.16 76.64 76.84 76.63
mDice 76.88 84.50 84.91 85.07 84.90
mRecall 76.26 88.00 89.89 90.48 89.00
mPrecision 79.06 83.41 84.37 83.63 83.62
Segformer mIoU 67.68 77.62 77.58 74.04 82.94
mDice 76.48 85.72 85.68 82.67 89.78
mRecall 80.61 92.99 91.87 86.55 95.15
mPrecision 79.50 84.27 83.68 79.84 90.87
Swin Trans. V2 mIoU 65.72 72.97 73.87 73.73 79.81
mDice 74.31 81.70 82.51 82.39 87.46
mRecall 74.41 83.58 85.14 85.00 91.43
mPrecision 77.76 80.03 80.34 80.16 86.71
@@
Figure 11: Segmentation metrics. In (a), the orange dots represent actual data points, and the gray dashed line indicates the estimated PR curve. A curve closer to upper right corner represents better. (b) illustrates the counterbalance between accuracy criterions (mDice) and positive pixel criterions (mRecall and mPrecision). The results demonstrate that our model can produce effective classification. Accuracy and positive detection performance reach the best balance point at $threshold\approx 0.3$.
@@ 4.4 Ablation of Distortion Parameter
We conducted ablation experiments on parameters $A$ and $\omega$ respectively as is shown in fig. 12. Overall, our model can achieve acceptable results within a large range of augmentation parameters, and the effects of adjacent parameters tend to be similar, which is manifested as a smoother surface in the performance distribution map. This feature ensures its ease of use, after all, researchers always tend to prefer a plug-and-play module rather than excessive parameter tuning.
@@
When $A$ or $\omega$ becomes too large, it will instead reduce the segmentation performance of the model. This is because excessive distortion will cause adjacent slice regions to be overly distorted, and the neural network cannot extract effective pixel features from these regions. Furthermore, larger $A,\omega$ parameters usually require larger $\delta$ to maintain the resolution of affine interpolation. The overall complexity of distortion can be approximated as $O(\delta^{2})$. If hardware constraints only allow the algorithm to operate with a smaller $\delta$, the augmented image may contain increased fragmentation. Noted that our similarity calculation is exactly designed to decrease the calculation cost during parameter search, and there do have correlations between the similarity metrics and the final segmentation metrics, as has been discussed in section 4.2.
@@ 5 Discussions
To our knowledge, our approach is one of the few instances within the Medical Imaging domain that employs image similarity metrics for neural network hyperparameter search [22]. Similar to the meta-data-driven Scan Table Removal technique, it offers significantly higher throughput compared to approaches relying solely on deep learning, while maintaining comparable accuracy. To implement these methods, researchers need to delve into a deeper understanding of medical imaging sequences. Given the remarkable efficacy of these non-deep-learning approaches, it’s encouraging to foresee an increase in similar algorithms tailored for applications in medical imaging.
@@ Acknowledgement
The work was supported in part by National Natural Science Foundation of China (Grant Nos. 61572325) and the Shanghai Key Technology Project (19DZ1208903).
I am profoundly grateful to my advisor (Prof. Chen) for generously providing the essential research environment, computational resources, and data resources that were crucial for the advancement of this study. I also extend my heartfelt thanks to my girlfriend (M. Eng. Chen) for her steadfast support and companionship, which greatly contributed to my perseverance during the most tough time.
@@
The implementation code used in this research is available online: https://github.com/MGAMZ/PSBPD. Due to limited resources, we currently only provide a preview version. A more comprehensive version will be updated after the article is accepted or published.

### v2

@@ 4.3 Similarity Guide Results
Our primary research goal is to develop a data augmentation technique that is interpretable, does not excessively alter samples, and adheres to clinical norms. The interpretability of the method cannot be assessed through end-to-end training alone. Initially, we employ the proposed image similarity algorithm to quantitatively evaluate whether the samples have been overly distorted by our method. A rapid decline in similarity at certain thresholds would indicate excessive distortion, suggesting lower interpretability as the augmented samples may not be readily comprehensible to clinical professionals.
The results show that with a variety of parameters employed for augmentation, there is a consistent similarity between the pre- and post-augmentation samples. This initially suggests that our proposed method does not encounter any uncontrollable divergences or singularities. Furthermore, there exists a parameter range that is relatively smooth, within which the method introduces minimal distortion to the images, making it more straightforward for the similarity detection algorithm to perform feature point matching. We observe that the number of successful SIFT pairings starts to plummet with $a\approx 3$$f\approx 1.5$ LABEL:fig:SimilarityGuide_Results. The Pearson correlation coefficient between model accuracy and two similarity metrics are $\rho_{SIRF-Acc}=-0.625$ and $\rho_{ORB-Acc}=-0.618$.
Given that our annotations are focused on gastric lesions, we have conducted an in-depth examination of the similarity levels across upper abdomen. Here, we aimed to assist in a more comprehensive determination of the optimal distortion intensity LABEL:fig:AdjacentSimilarityDistribution. The similarity criterion is described in Eq. 14. A higher $N_{\triangle}$ indicates a greater complexity of textures in the respective axial position, enabling the extraction of a larger number of feature points.
@@
While similarity detection is initially employed to ascertain a general range for the augmentation parameters, Section 4.5 involved training each parameter from scratch. This approach is taken to validate the reliability of the ideal parameter range that is more efficiently derived through similarity detection.
4.4 Segmentation Evaluations
We examine our method’s effectiveness the latest segmentation models, MedNext[30] and SwinUMamba[16], combined with the large dataset provided by SA-Med2D[40]. The results are presented in Tables 4 and 5.
Table 4: Accuracy on the large CT-ORG[29] dataset. Our method is able to help the latest neural networks to achieve higher accuracy with limited training samples, specially useful in medical scenarios. This dataset is a typical representative of CT modality.
@@ 4.4 Segmentation Evaluations
MedNext Rotation ±$\pm 90^{\circ}$ Dice 96.39 65.53 3.21 19.32 66.26 86.13 56.14
@@
SwinUMamba Rotation ±$\pm 90^{\circ}$ Dice 96.85 71.44 1.39 21.38 71.84 92.3 59.21
@@
Table 5: The similar experiment with Table 4, but is on the BraTS[22] dataset, and trained with SwinUMamba. This dataset is a typical representative of MR modality.
@@
BraTS 2020 Rotation ±$\pm 90^{\circ}$ Dice 99.06 37.69 13.75 26.93 44.36
@@
BraTS 2021 Rotation ±$\pm 90^{\circ}$ Dice 99.03 35.24 1.53 30.76 41.64
@@
In Table 6, we evaluated our method on several traditional and widely-used neural networks, training on our private gastric cancer dataset. The dataset allows us to examine the effectiveness of the proposed Metadata-Driven Scan Table Removal and Distortion Augmentation. The results show an improvement on segmentation accuracy across multiple neural-network-based segmentation frameworks, which are selected from the most representative ones in recent computer vision researches. Most frameworks can steady gain higher value on major metrics without the extra samples. Our method has demonstrated a consistent enhancement in Dice scores across various experiments, excluding the Poolformer model. The improvement is evident even when distortion is utilized without the Table Removal feature. However, the incremental gain in accuracy when Table Removal is activated is relatively minor, suggesting its impact may not be as significant.
@@
Family Model Criterion w/o aug1 rotate ±$\pm 45^{\circ}$ rotate ±$\pm 90^{\circ}$ rotate ±$\pm 180^{\circ}$ w/o Table Removal Ours
Conv-Based Resnet50 mIoU 72.52 78.32 79.49 78.58 81.27 82.57
mDice 81.28 86.28 87.20 86.48 88.56 89.52
mRecall 82.39 87.34 88.67 91.53 88.90 91.37
mPrecision 83.81 87.04 87.80 85.87 89.07 89.05
ConvNext mIoU 63.47 71.99 75.2 75.00 77.46 79.64
mDice 71.72 80.77 83.75 83.51 85.58 87.32
mRecall 77.36 82.26 84.36 86.95 88.68 90.37
mPrecision 75.54 82.57 83.75 82.83 85.94 87.16
SegNeXt mIoU 64.02 72.46 73.04 74.16 76.37 76.61
mDice 72.37 81.23 81.77 82.76 84.67 84.88
mRecall 70.31 83.81 86.20 85.98 87.06 89.39
mPrecision 77.46 80.73 82.54 83.60 84.34 85.57
Trans-Based MAE mIoU 67.81 71.73 70.00 68.90 75.92 75.04
mDice 76.63 80.56 78.89 77.81 84.30 83.55
mRecall 85.79 85.80 82.25 84.59 87.84 87.09
mPrecision 78.27 77.44 78.66 73.90 81.86 82.70
Poolformer mIoU 68.08 76.16 76.64 76.84 78.60 76.63
mDice 76.88 84.50 84.91 85.07 86.50 84.90
mRecall 76.26 88.00 89.89 90.48 88.98 89.00
mPrecision 79.06 83.41 84.37 83.63 84.33 83.62
Segformer mIoU 67.68 77.62 77.58 74.04 81.19 82.94
mDice 76.48 85.72 85.68 82.67 88.51 89.78
mRecall 80.61 92.99 91.87 86.55 93.40 95.15
mPrecision 79.50 84.27 83.68 79.84 85.14 90.87
Swin Trans. V2 mIoU 65.72 72.97 73.87 73.73 76.34 79.81
mDice 74.31 81.70 82.51 82.39 84.65 87.46
mRecall 74.41 83.58 85.14 85.00 87.96 91.43
mPrecision 77.76 80.03 80.34 80.16 83.48 86.71
@@
In the context of MAE, the exclusion of the Table Removal process yields superior accuracy. This is predominantly due to the masked reconstruction phase inherent in MAE’s learning mechanism, which affords equal consideration to all pixels within the slice, thereby not treating the CT Scan Table as noise. The removal of the Table leads to a reduction in image information entropy, which simplifies the learning task for MAE and diminishes the learning space. As a result, the model’s ultimate accuracy is compromised.
Fig. 11 illustrate the Precision-recall Curve. Due to limited computational resources, we focused on calculating the most dynamically changing segmentgain higher value on of the PR curve.
Figure 11: Segmentation metrics. In the left subfigure, the orange dots represent actual data points, and the gray dashed line indicates the estimated PR curve. A curve closer to upper right corner represents better. The right subfigure illustrates the counterbalance between accuracy criterions (mDice) and positive pixel criterions (mRecall and mPrecision). The results demonstrate that our model can produce effective classification.
@@ 4.5 Ablation of Distortion Parameter
We conducted ablation experiments on parameters $A$ and $\omega$ respectively as is shown in Fig. 12. Overall, our model can achieve acceptable results within a large range of augmentation parameters, and the effects of adjacent parameters tend to be similar, which is manifested as a smoother surface in the accuracy distribution map. This feature ensures its ease of use, after all, researchers always tend to prefer a plug-and-play module rather than excessive parameter tuning.
@@
When $A$ or $\omega$ becomes too large, it will instead reduce the segmentation accuracy of the model. This is because excessive distortion will cause adjacent slice regions to be overly distorted, and the neural network cannot extract effective pixel features from these regions. Furthermore, larger $A,\omega$ parameters usually require larger $\delta$ to maintain the resolution of affine interpolation. The overall complexity of distortion can be approximated as $O(\delta^{2})$. If hardware constraints only allow the algorithm to operate with a smaller $\delta$, the augmented image may contain increased fragmentation.
When comparing the results to Section 4.3, there appears to be a correlation between the number of matching points identified by the similarity algorithm and the accuracy of the end-to-end prediction. The neural network’s accuracy starts to falter rather than improve when the augmentation intensity with $A>7,\omega>3$, and the number of successful pairings during similarity calculation starts to plummet with $a\approx 3$$f\approx 1.5$. Noted that during training, we use Eq. 2 to randomly determine $A,\omega$ for each training batch, while use constant value during similarity calculation. Consequently, the mean of the sampling distribution for random parameters during the training phase can serve as a benchmark for comparison with the parameter values applied in the similarity detection process. Based on these observations, it is evident that there is a similarity between the distributions in question. The similarity detection and the neural network’s predictions for slices exhibit a rapid decline in segmentation accuracy under similar augmentation intensities. This implies that the proposed similarity detection algorithm has the potential to accurately estimate the optimal parameter range without the prerequisite of actual neural network training, thereby substantially reducing the computational expense associated with parameter adjustment.
4.6 Computation Complexity
In line with the theoretical analysis provided in Section 3.5, our implementation identifies the determination of the affine mapping matrix and the mapping sampling process as the primary computational bottlenecks, which are correlated with the quantity of control points and source pixels, respectively. We conducted ablation experiments on these two factors, and the findings are presented in Fig. 13.
Figure 13: Computational Overhead Analysis. Distinct shapes are utilized to denote varying sizes $n$ of the source pixel matrices. Different colors are assigned to indicate the selection of post-sampling implementation options. Additionally, the size of the shapes corresponds to the relative memory footprint. In summary, the efficiency of Scipy is inferior, whereas cv2 consistently demonstrates millisecond-level responsiveness across all configurations.
The experimental findings indicate a positive correlation between the number of control points (Grid Density) and processing time. Additionally, an increase in the resolution of the source matrix results in higher memory consumption. When employing OpenCV for mapping sampling instead of Scipy[37], the study observed substantially reduced processing times and lower memory usage. As the computation method for the mapping matrix remained unchanged, utilizing the same piecewise affine mapping technique with Delaunay triangulation, it is inferred that the primary time overhead is attributed to the resampling of the source matrix.
The aforementioned data is obtained through the process of independent data augmentation execution. In order to better stick to the practical application scenarios, we conducted a further analysis to ascertain whether data preprocessing acts as a bottleneck in the computational speed of neural networks during model training, thereby potentially causing periods of GPU idleness. The results are presented in Table 7. The GPU exhibits minimal idleness during the training process. This is attributed to the CPU in completing the preprocessing of subsequent batch of data in advance of the completion of the current batch’s neural network computations, which further indicating the efficiency of the proposed method.
Table 7: Computational Overhead Analysis. The table lists the average time taken for each step in the data augmentation process. The time unit is in milliseconds. The results indicate that the data augmentation process is not a bottleneck in the training process.
Model Batch Time GPU Idle
ConvNeXt 200 2.2
MAE 271 2.0
Poolformer 142 3.5
ResNet50 259 1.7
Segformer 239 2.6
SegNeXt 123 3.2
SwinT. V2 194 3.9
@@ 5 Discussions
To our knowledge, our approach is one of the few instances within the Medical Imaging domain that employs image similarity metrics for neural network hyperparameter search [24]. Similar to the meta-data-driven Scan Table Removal technique, it offers significantly higher throughput compared to approaches relying solely on deep learning, while maintaining comparable accuracy. To implement these methods, researchers need to delve into a deeper understanding of medical imaging sequences. Our research suggests that while the approach to determining control points for segmented affine mapping is intuitive, this approach may inadvertently constrain the diversity of outcomes. Merely designating the area adjacent to the spine as the focal point for distortion may not yield optimal results. Given the current capability in the academic community to automate the localization of numerous human organs, tissues, and structures, applying a regional-level segmented affine mapping individually to each target instance based on these localization results could potentially enhance the robustness of the outcomes and increase the variety of the samples generated.
In the context of X-Ray imaging, it is common to obtain projections of slices exclusively on the coronal plane. This limitation implies that the method introduced in this research is not readily applicable to Chest X-Rays (CxR). Given the greater prevalence of X-Rays compared to CT scans, a technique that circumvents the need for slice modeling would enhance the scope of augmentation.
Our objective is to capitalise on the distinct attributes of medical 3D Volume imagery in contrast to traditional visual imaging or point cloud 3D imagery. We intend to develop a highly persuasive augmentation approach that harnesses the inherent benefits of medical 3D Volume images. In the realm of healthcare, it is imperative that not only neural networks but every component of the implementation garners sufficient credibility within clinical contexts. This level of trust is a fundamental requirement for the effective fusion of artificial intelligence with medical practices.
@@ Acknowledgement
The work is supported in part by National Natural Science Foundation of China (Grant Nos. 61572325) and the Shanghai Key Technology Project (19DZ1208903).
I am profoundly grateful to my advisor (Prof. Chen) for generously providing the essential research environment, computational resources, and data resources that are crucial for the advancement of this study. I also extend my heartfelt thanks to my girlfriend (M. Eng. Chen) for her steadfast support and companionship, which greatly contributed to my perseverance during the most tough time.
@@
The implementation code used in this research is available online: https://github.com/MGAMZ/PSBPD.
[…]